# Subtype-Aware Registration of Longitudinal Electronic Health Records

### v1

@@ 2.1 Step 3: Finalization
6:Output final shift amounts $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{S}_{k}^{(t_{\text{final}})}\}$
3 Simulations
@@
In this section, we conduct simulation studies across six scenarios to evaluate the effectiveness of our proposed algorithm. These scenarios covered a variety of different longitudinal data structures, including different shapes, different means, and their combinations. The objective is to assess the proposed algorithm’s ability to correctly recover the true temporal positions of these longitudinal data under various conditions. Next, we will describe the common setup, then discuss specific designs, and finally discuss the simulation results.
Common Simulation Setup.
@@
Algorithm Parameter Settings: We fix 2 internal knots at 8 and 13. The potential shift positions $\Lambda$ were set to $\{0,1,2,3,4\}$. The range of cluster numbers was set to $\{2,3,\ldots,8\}$, i.e., $M=8$. The selection proportion $\alpha$ was set to 0.95, the threshold $\tau$ was set to 0.45, and the maximum iteration numbers were set to $I_{\text{small}}=1$ and $I_{\text{large}}=10$. We employ $k$-medoids as the clustering method employed given its robustness to outliers [35].
@@
This scenario involves two types of trajectories that are similar in shape but have significant differences in mean, defined as follows:
$\begin{split}\text{Group 1 (500 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (500 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In this setting, the trajectories from Group 1 had a higher mean level compared to Group 2, while the overall shape of their trajectories, dictated by the sinusoidal base sequence, remained similar. This setup allows us to evaluate the algorithm’s ability to distinguish between groups with different mean levels but similar temporal patterns.
@@
This scenario involves two types of trajectories with distinct shape differences but similar overall means.
$\begin{split}\text{Group 1 (500 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=17+2\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (500 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=2t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
While both groups have similar overall means, Group 1 follows a sinusoidal pattern, whereas Group 2 exhibits a combination of a sinusoidal pattern and a linear trend.
@@
This scenario involves three types of trajectories defined as follows:
$\begin{split}\text{Group 1 (300 subjects): }\quad&Y_{i}^{(1)}(t_{ij})=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (400 subjects): }\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 3 (300 subjects): }\quad&Y_{i}^{(3)}(t_{ij})=16+0.5t_{ij}+3\sin(0.9(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In summary, Groups 1 and 2 have similar shapes but different means, while Group 3 has a different shape that overlaps with the first two groups. This scenario helps evaluate the algorithm’s ability to distinguish between groups with both shape and mean differences.
@@
This scenario involves four types of trajectories with shape differences and overlaps:
$\begin{split}\text{Group 1 (250 subjects):}\quad&Y_{i}^{(1)}(t_{ij})=24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 2 (250 subjects):}\quad&Y_{i}^{(2)}(t_{ij})=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij};\\ \text{Group 3 (250 subjects):}\quad&Y_{i}^{(3)}(t_{ij})=38-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij};\\ \text{Group 4 (250 subjects):}\quad&Y_{i}^{(4)}(t_{ij})=21+0.9t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.\end{split}$
In this scenario, Groups 1 and 2 have similar shapes but different means, while Groups 3 and 4 exhibit different shapes that overlap with each other and the first two groups.
@@
This scenario includes 4 groups of trajectories and each group includes 250 subjects. The trajectories of each group is generated as follows:
$\begin{split}\text{Group 1 (250 subjects):}\quad Y_{i}^{(1)}(t_{ij})=&24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}^{(1)};\\ \text{Group 2 (250 subjects):}\quad Y_{i}^{(2)}(t_{ij})=&17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}^{(2)};\\ \text{Group 3 (250 subjects):}\quad Y_{i}^{(3)}(t_{ij})=&28-0.5t_{ij}+4\sin(0.9(t+3))+\epsilon_{ij}^{(3)};\\ \text{Group 4 (250 subjects):}\quad Y_{i}^{(4)}(t_{ij})=&21-0.5t_{ij}+4\sin(0.9(t+3))+\epsilon_{ij}^{(4)}.\end{split}$
In this scenario, the trajectories in Groups 1 and 2 share one shape with different means, while Groups 3 and 4 share another shape with different means. This scenario helps evaluate the algorithm’s ability to distinguish between overlapping groups with similar shapes but different mean levels.

### v2

@@ 3 Method for Subtype-Aware Registration of EHR
6:Output final shift amounts $\{\delta_{i}^{*}\mid i=1,2,\ldots,N\}$ and clusters $\{\mathcal{S}_{k}^{(h_{\text{final}})}\}$
In this section, we apply the proposed registration method to longitudinal SCr trajectories from the MIMIC-IV ICU AKI cohort described in Section 2.
4.1 Experimental Setup
Dataset and Cohort Selection
We use the MIMIC-IV AKI ICU cohort defined in Section 2, consisting of ICU stays with longitudinal serum creatinine (SCr) trajectories over the first 21 days after ICU admission.
Data Preprocessing and Model Selection
We perform unsupervised learning on the data. Specifically, we transform the original longitudinal data as detailed in Section 3, and used the estimated coefficients of the spline basis functions to represent each ICU stay. We applied $k$-means clustering macqueen1967some and $k$-medoids clustering rdusseeun1987clustering on these representations. The quality of the resulting clusters was evaluated by computing the average silhouette coefficient rousseeuw1987silhouettes .
For clinical evaluation, we labeled each ICU stay based on the AKI severity stages defined below Khwaja2012KDIGOCP ; Zhu2024FunctionalPS :
  1. 1.
    Stage 1: Serum creatinine (SCr) level rises to $1.5\times$ to $1.9\times$ the baseline value within seven days since time 0.
  2. 2.
    Stage 2: SCr level rises to $2.0\times$ to $2.9\times$ the baseline value within seven days since time 0.
  3. 3.
    Stage 3: SCr level rises to $3\times$ or greater than the baseline value within seven days, or the maximum SCr level over two days is $\geq 4.0\;\text{mg/dL}$. We further divide Stage 3 into two subtypes Zhu2024FunctionalPS :
    1. (a)
      Subtype 1: SCr level $\geq 4.0\;\text{mg/dL}$ within 48 hours since time 0.
    2. (b)
      Subtype 2: SCr level rises to $3\times$ or greater than the baseline value within seven days since time 0 and does not meet the criteria for Subtype 1.
The baseline SCr value was computed using the Modification of Diet in Renal Disease (MDRD) equation Levey1999AMA .
We also perform supervised learning on the data. We specifically organize the original longitudinal EHR data into an $N\times T$ array, where $N$ is the number of ICU stays and $T$ represents the number of days since ICU admission. We performed data normalization and missing data imputation following the methods described in Jiang2023SoftPF . We computed summary statistics (maximum, minimum, mean, standard deviation, and skewness) over 3-day and 4-day windows and performed severe AKI prediction using logistic regression, based on the methods described in Lipton2016LearningTD . We defined severe AKI as observing a $3\times$ increase in creatinine level compared to the baseline value or observing any creatinine level $\geq 4.0\;\text{mg/dL}$ during the whole observed time interval of 21 days Khwaja2012KDIGOCP . For the baseline method in jiang2023timeline , we set the potential time shifts to within seven days and the maximum iteration number to 10.
4.2 Unsupervised Learning – Phenotyping
We first evaluate the impact of timeline registration on clustering AKI ICU stays using $k$-medoids and $k$-means. The average silhouette coefficients for different numbers of clusters $K\in[2,8]$ are presented in Table 1. Notably, when $K=2$ and $K=3$, the clustering results yield significantly higher average silhouette coefficients, indicating more distinct clustering results. Additionally, we observe that both $k$-medoids and $k$-means show substantial increases in the average silhouette coefficients after applying the proposed timeline registration method. Given the heterogeneity of AKI, we expect the ICU stays to cluster according to different subtypes of the condition. The numerical results demonstrate the effectiveness of our timeline registration method in aligning the timelines of patients within the same subtype. Furthermore, our proposed registration method clearly outperforms the baseline registration method, achieving more distinct subtype separation for most choices of $K$.
$k$-medoids $k$-means
Observed Data Registered Data Registered Data (baseline) Observed Data Registered Data Registered Data (baseline)
KK=2 0.431 0.481 0.266 0.584 0.621 0.518
KK=3 0.226 0.261 0.326 0.523 0.557 0.497
KK=4 0.249 0.276 0.276 0.325 0.328 0.479
KK=5 0.208 0.278 0.086 0.350 0.380 0.307
KK=6 0.208 0.282 0.075 0.352 0.394 0.330
KK=7 0.111 0.164 0.075 0.346 0.239 0.328
KK=8 0.108 0.156 0.025 0.284 0.220 0.296
Table 1: Comparison of average silhouette coefficients for original and registered data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering.
In computational phenotyping research, the number of subtypes $K$ is often unknown and must be determined. Therefore, we consider a scenario in which the number of clusters falls within a pre-specified range, and we apply the proposed timeline registration method to determine this number in a data-driven manner. In Table 2, we present results for two scenarios: one requiring a smaller number of clusters and the other requiring a larger number, as potentially suggested by auxiliary information. In both cases, we observe that the maximum average silhouette coefficient increased after timeline registration, indicating that the timeline registration method provides consistent improvement in downstream clustering performance.
$k$-medoids $k$-means
Observed Data Registered Data Observed Data Registered Data
$K\in[2,4]$ 0.431 0.482 0.584 0.621
$K\in[5,8]$ 0.208 0.284 0.352 0.432
Table 2: Comparison of the maximum average silhouette coefficients obtained within the cluster ranges $K\in[2,4]$ and $K\in[5,8]$ for the original data and the registered data after applying the proposed timeline registration method, using $k$-medoids and $k$-means clustering.
4.2.1 Phenotype Interpretation and Evaluation
Next, we further explored and evaluated the phenotyping results before and after temporal registration. From a precision-medicine perspective, clinically meaningful phenotypes are not only associated with distinct biology and outcomes, but also need to be identifiable early enough in the disease course to inform prevention, timely interventions, and resource allocation matthay2020phenotypes ; wick2021promises ; mcdonald2019treatable . For clarity, we denote phenotypes obtained from the raw (unregistered) trajectories by $S_{k}^{\mathrm{raw}}$ and those obtained from the time-registered trajectories by $S_{k}^{\mathrm{rec}}$.
Based on the unsupervised clustering results described above, we focused on the clustering solution with $K=6$, which achieved the largest average silhouette coefficient among the more fine-grained clusterings considered (i.e., for $K\geq 3$). In this analysis, we divided the 21-day observation window into an early period (days 1–10), a mid period (days 11–15), and a late period (days 16–21), and defined a SCr-based outcome as the maximum creatinine value observed in the late period. We then evaluated how reliably phenotypes could be identified in the early and intermediate windows. As shown in Figure 2, for the raw data, the two phenotypes that can be identified in the early and mid windows are $S_{4}^{\mathrm{raw}}$ and $S_{6}^{\mathrm{raw}}$, and only $S_{6}^{\mathrm{raw}}$ remains clearly associated with a distinct SCr-based outcome. In contrast, for the registered data the three phenotypes that are most easily identifiable in the early and mid windows are $S_{3}^{\mathrm{rec}}$, $S_{5}^{\mathrm{rec}}$, and $S_{6}^{\mathrm{rec}}$, and all three exhibit well-separated distributions of this SCr-based outcome.
Figure 2: Early and mid SCr trajectories and SCr-based outcome by subtype. For the early and mid periods (definitions given in the Methods), the curves display the mean SCr over time and the shaded bands represent mean $\pm 0.5$ SD within each subtype. The right-hand panels summarise the SCr-based outcome, defined as each patient’s peak SCr, i.e., the maximum SCr value observed in the late period.
These findings suggest that phenotyping based on the registered data yields phenotypes that are more clinically meaningful and potentially more actionable: once a longitudinal phenotype can be recognized early, we gain phenotype-specific information about the subsequent course, which in turn can guide targeted treatment and nursing strategies and facilitate the implementation of personalized care.
To further evaluate these phenotypes, we used diagnosis text variables to characterize the three time-registered phenotypes. We first excluded diagnosis entries that explicitly referred to “acute kidney injury” or “acute kidney failure”, so that AKI-defining terms themselves were not used in the downstream characterization. Among the remaining diagnoses, we then performed a focused semantic harmonization for a set of predefined, clinically important diagnosis terms by grouping them into several categories, including hyperkalemia, hyponatremia and/or hyposmolality, acute respiratory failure, the sepsis spectrum, essential (primary) hypertension, hyperlipidemia, atrial fibrillation, ascites, urinary tract infection, and chronic kidney disease, mapping all synonymous textual descriptions of each of these conditions to the corresponding category. Using these harmonized categories, we summarized the prevalence of key diagnoses (proportion of patients with each diagnosis) across $S_{3}^{\mathrm{rec}}$, $S_{5}^{\mathrm{rec}}$, and $S_{6}^{\mathrm{rec}}$ (Figure 3), which provided a concise overview of their diagnostic profiles. This led to the following clinical interpretation of these three AKI phenotypes:
Figure 3: Top 20 diagnoses by patient proportion for the registered subtypes $S_{3}^{\mathrm{rec}}$, $S_{5}^{\mathrm{rec}}$, and $S_{6}^{\mathrm{rec}}$.
Phenotype $S_{3}^{\mathrm{rec}}$ (a sepsis-dominant AKI phenotype) is characterized by a high burden of sepsis-spectrum diagnoses together with acute respiratory failure, metabolic acidosis, and moderate electrolyte disturbances (hyponatremia/hyposmolality). Its clinical profile is broadly similar to the clinical picture of sepsis-associated AKI described in the recent ADQI consensus report by Zarbock et al. zarbock2023sepsis .
Phenotype $S_{5}^{\mathrm{rec}}$ (a metabolic-electrolyte derangement AKI phenotype) is characterized by a high prevalence of acute respiratory failure, acidosis, hyponatremia/hyposmolality, hyperkalemia, and acute post-hemorrhagic anemia. This clinical profile shares key features with critically ill patients in whom AKI is accompanied by severe metabolic and electrolyte complications, regarded as classic indications for urgent kidney replacement therapy (rahman2012acute, ).
Phenotype $S_{6}^{\mathrm{rec}}$ (a cardiopulmonary failure with multi-organ dysfunction AKI phenotype) shows frequent acute respiratory failure and relatively high prevalence of atrial fibrillation, sepsis-spectrum diagnoses, metabolic acidosis, chronic kidney disease, and other organ dysfunctions. Its clinical profile is close to AKI occurring with acute cardiopulmonary failure (for example, acute decompensated heart failure as a typical cardiorenal syndrome type 1 scenario) and with acute respiratory failure/ARDS, where lung-kidney cross-talk has been linked to organ dysfunction and adverse outcomes (rangaswami2019cardiorenal, ; HusainSyed2016LungKidney, ).
Taken together, these similarities indicate that temporal registration substantially improves phenotyping: after registration, the clustering recovers clinically meaningful AKI patterns that can be used as interpretable building blocks in subsequent analyses.
4.3 Clinical Evaluation
Next, we incorporate clinical information, specifically the AKI severity stages, into the evaluation of timeline registration performance. Recall that the definition of AKI severity stages, as outlined in Section 4.1, relies on the starting time point. This implies that the AKI severity stage of an individual subject may change after applying timeline registration. Therefore, for both the observed data (without timeline registration) and the data after applying the proposed timeline registration procedure, we label each subject as Stage 1 (S1), Stage 2 (S2), Stage 3 Subtype 1 (S3-1), or Stage 3 Subtype 2 (S3-2). We then group both the observed and registered datasets into five different configurations, as described in the first column of Table 3, and evaluate the average silhouette coefficient for each configuration. The results for these five different AKI sub-category divisions are presented in Table 3. Consistently, the data processed through timeline registration demonstrate an increased average silhouette coefficient, indicating improved data aggregation within each AKI sub-category. The baseline registration method also shows good performance in the first four configurations. However, it underperforms compared to the observed data in the fifth configuration. In contrast, the proposed registration method consistently outperforms the observed data across all configurations, demonstrating the robustness of our approach.
Observed Data Registered Data (baseline) Registered Data ($k$-means) Registered Data ($k$-medoids)
4-Groups (S1, S2, S3-1, S3-2) -0.049 -0.038 -0.039 -0.041
3-Groups (S1, S2, (S3-1, S3-2)) -0.057 -0.040 -0.046 -0.048
3-Groups ((S1, S2), S3-1, S3-2) 0.286 0.335 0.326 0.322
2-Groups ((S1, S2), (S3-1, S3-2)) 0.321 0.341 0.342 0.338
2-Groups ((S1, S2, S3-1), S3-2) 0.527 0.503 0.625 0.663
Table 3: Comparison of observed data and registered data in the clinical setting using the new criteria for defining creatinine baseline level. For each AKI sub-category configuration, the reported values are the average silhouette coefficients. Larger values indicate better aggregation and clearer separation of subjects within the specified AKI sub-categories.
4.4 Supervised Learning – Severe AKI Prediction
Observed Data Registered Data (baseline) Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8047 [0.7801,0.8275] 0.7441 [0.7202,0.7687] 0.8559 [0.8357,0.8734] 0.8467 [0.8243,0.8679]
Precision 0.8016 [0.7765,0.827] 0.8326 [0.7621,0.8637] 0.8235 [0.801,0.8454] 0.8181 [0.7942,0.841]
Recall 0.7139 [0.6944,0.7366] 0.5114 [0.5056,0.5171] 0.7546 [0.7328,0.7752] 0.7454 [0.7241,0.7654]
Accuracy 0.8167 [0.8031,0.8312] 0.7268 [0.7234,0.7303] 0.8385 [0.8247,0.8514] 0.8333 [0.8193,0.8473]
AUPRC 0.8022 [0.7803,0.8227] 0.7393 [0.7115,0.7635] 0.8446 [0.824,0.8618] 0.8355 [0.8139,0.8566]
Observed Data Registered Data (baseline) Registered Data ($k$-medoids) Registered Data ($k$-means)
AUROC 0.8274 [0.8033,0.8497] 0.7752 [0.7493,0.7991] 0.8788 [0.8605,0.895] 0.8687 [0.8475,0.8875]
Precision 0.8143 [0.7896,0.84] 0.7357 [0.7123,0.7607] 0.8351 [0.8134,0.855] 0.8353 [0.8146,0.8578]
Recall 0.733 [0.7121,0.7531] 0.7127 [0.6897,0.7365] 0.7721 [0.7515,0.7929] 0.7694 [0.7495,0.7896]
Accuracy 0.8276 [0.8138,0.8412] 0.7862 [0.7682,0.8032] 0.8488 [0.8357,0.8617] 0.848 [0.8343,0.861]
AUPRC 0.8257 [0.8043,0.8472] 0.7597 [0.7329,0.7849] 0.8689 [0.8509,0.8846] 0.8595 [0.8394,0.8775]
From both tables, we observe that using data after registration results in significant increases in the AUROC (area under the receiver operating characteristic curve) and AUPRC (area under the precision-recall curve) values. Accuracy values also slightly increase. This comparison indicates that timeline registration is influential in enhancing downstream clinical analysis and can be considered an essential preprocessing step. We also observe that the baseline EHR registration method yields results worse than using data without registration from both tables. This can be attributed to the reason that the baseline method relies on the assumption that all patients in the cohort share the same disease progression timeline. Since AKI patients may undergo distinct disease progression patterns, enforcing patient data to follow a shared trend may result in inaccurate timeline alignment between patients and thus degrading the prediction performance.
4.4.1 Subgroup Analysis
We conducted subgroup analysis on groups by age levels Xu2021SystematicRA (group of patients that are below 60 years of old and group of patients over 60 years of old) and gender. We present the performance of the prediction model on observed data and data recovered by our proposed method in each subgroup shown in Table 6 and Table 7. The experiments were conducted 500 times using 4-day data with the mean values and 95% confidence intervals reported.
We observe from the tables that the prediction model yielded higher performance using registered data obtained from the proposed method than using the original data in both gender groups. Likewise, we note a higher performance of the model with the use of proposed data registration in both age groups. Additionally, we show the prediction model yielded comparable performance in all subgroups.
Overall, early warning of AKI across diverse clinical settings is clinically important song2020cross ; zhang2025development , and our results suggest that the proposed temporal registration method can improve the performance of such supervised prediction models.
AUROC Precision Recall Accuracy AUPRC
Male Observed Data 0.8357 [0.8088,0.8625] 0.7974 [0.7659,0.8271] 0.7459 [0.7174,0.7733] 0.8318 [0.8124,0.8511] 0.833 [0.8055,0.8583]
Registered Data (baseline) 0.8018 [0.7692,0.8297] 0.7573 [0.7297,0.7864] 0.7507 [0.7235,0.7776] 0.8104 [0.7893,0.8305] 0.7791 [0.7421,0.8108]
Registered Data ($k$-medoids) 0.8983 [0.8755,0.918] 0.839 [0.8109,0.8638] 0.8083 [0.7808,0.8319] 0.867 [0.8482,0.8855] 0.8865 [0.8642,0.9066]
Registered Data ($k$-means) 0.8842 [0.8635,0.9057] 0.8295 [0.8028,0.857] 0.7981 [0.7705,0.8217] 0.8599 [0.8415,0.8797] 0.8739 [0.8535,0.8941]
Female Observed Data 0.8288 [0.7935,0.8631] 0.8564 [0.8259,0.8836] 0.7178 [0.6865,0.7496] 0.821 [0.7975,0.8463] 0.8327 [0.7992,0.8639]
Registered Data (baseline) 0.7448 [0.7056,0.7862] 0.7007 [0.6649,0.7415] 0.6594 [0.6268,0.6951] 0.7472 [0.7167,0.774] 0.74 [0.7043,0.7801]
Registered Data ($k$-medoids) 0.8641 [0.8332,0.8906] 0.8386 [0.8074,0.8679] 0.723 [0.6917,0.7531] 0.8196 [0.796,0.8428] 0.858 [0.8271,0.8835]
Registered Data ($k$-means) 0.8594 [0.8249,0.8921] 0.8599 [0.8274,0.8868] 0.7317 [0.7006,0.7638] 0.8288 [0.8067,0.8536] 0.8558 [0.8244,0.8867]
Table 6: Subgroup analysis on gender. Metrics are presented using mean values with 95% confidence intervals.
AUROC Precision Recall Accuracy AUPRC
Age <60<60 Observed Data 0.8478 [0.8139,0.8838] 0.818 [0.7766,0.8576] 0.7662 [0.7332,0.8038] 0.8483 [0.8241,0.875] 0.8479 [0.8159,0.8795]
Registered Data (baseline) 0.8099 [0.7724,0.8444] 0.7658 [0.7305,0.8012] 0.748 [0.7103,0.7859] 0.8185 [0.7924,0.8439] 0.7871 [0.7428,0.825]
Registered Data ($k$-medoids) 0.9007 [0.8731,0.9265] 0.8359 [0.8001,0.8687] 0.7955 [0.761,0.832] 0.8641 [0.8402,0.8868] 0.8888 [0.8623,0.9147]
Registered Data ($k$-means) 0.8945 [0.866,0.9219] 0.8429 [0.8102,0.8747] 0.8123 [0.7795,0.845] 0.8718 [0.848,0.896] 0.8875 [0.8604,0.9145]
Age $\geq 60$ Observed Data 0.8163 [0.7883,0.8432] 0.8137 [0.7866,0.8403] 0.7159 [0.6898,0.7433] 0.8153 [0.7974,0.8341] 0.8139 [0.7857,0.8391]
Registered Data (baseline) 0.7551 [0.7241,0.7841] 0.7178 [0.6869,0.7484] 0.6932 [0.6657,0.7221] 0.7669 [0.7433,0.789] 0.7446 [0.7117,0.7754]
Registered Data ($k$-medoids) 0.8664 [0.8424,0.8871] 0.8354 [0.8079,0.8613] 0.76 [0.7337,0.7861] 0.8397 [0.8213,0.8593] 0.8579 [0.8348,0.8793]
Registered Data ($k$-means) 0.8556 [0.8262,0.8808] 0.8327 [0.8058,0.859] 0.7472 [0.7201,0.7739] 0.8337 [0.8153,0.8529] 0.8451 [0.8177,0.8692]
Table 7: Subgroup analysis on age levels. Metrics are presented using mean values with 95% confidence intervals.
5 Discussions and Conclusions
In this paper, we demonstrate the necessity of performing timeline registration for longitudinal EHR data, aligning patient records to reflect the intrinsic disease progression trends. To evaluate the feasibility of the proposed registration algorithm for recovering ground truth longitudinal data, we performed simulations with eight different scenarios. The results indicate that the proposed method can effectively recover observed data to ground truth with low time complexity. To demonstrate the impact of registration on downstream clinical data analyses, we perform phenotyping, clinical evaluation and severe AKI prediction and compare their results using data with or without timeline registration. The results suggest that registration can effectively improve the performance of downstream clinical data analysis. We also conclude that the registration algorithm is of high stability, reflected by the minimal 95% confidence intervals given 500 times repetition of experiments. Moreover, the proposed framework is flexible and readily extensible. Both the dimension-reduction and clustering steps can be substituted with alternative methods according to users’ preferences and practical needs, provided that the chosen dimension-reduction method preserves separability among distinct trajectory patterns and the clustering method is capable of yielding a meaningful grouping structure. This flexibility is empirically illustrated in Supplementary Section A through ablation studies that examine representative alternatives for these components. On the other hand, our method also has some limitations. For example, in diseases with rare trajectory subtypes, the accuracy of alignment for these rare subtypes may be less accurate due to the potential overshadowing effect from other more dominant subtypes. Registration on continuous space and evaluation of registration on other critical medical events will be investigated in our future work. EHR timeline registration is an emerging field and our work provides insight for improving clinical analysis using EHR data.
Supplementary Materials for “Subtype-Aware Registration of Longitudinal Electronic Health Records”
Appendix A Simulations
@@
In this section, we conduct simulation studies across eight scenarios to evaluate the ability of the proposed algorithm to recover the true temporal positions of longitudinal trajectories under diverse shapes, mean structures, and progression speeds. We first describe the common setup, then the specific designs, and finally the results.
A.1 Common Simulation Setup
@@
Algorithm Parameter Settings: We fix 2 internal knots at 8 and 13. The potential shift positions $\Lambda$ were set to $\{0,1,2,3,4\}$. The range of cluster numbers was set to $\{2,3,\ldots,8\}$, i.e., $M=8$. The selection proportion $\alpha$ was set to 0.95, and the threshold $\tau$ was set to 0.45. We employ $k$-medoids as the clustering method employed given its robustness to outliers rdusseeun1987clustering .
A.2 Scenario Design
Eight scenarios are constructed to represent different combinations of trajectory characteristics:
  • Scenarios 1–2: Distinguish trajectories based on either different means or different shapes.
  • Scenarios 3–4: Incorporate both shape and mean differences with overlapping patterns and increased cluster complexity.
  • Scenarios 5–6: Add intra-shape variation in mean levels and increased inter-group similarity.
  • Scenarios 7–8: Introduce within-group progression speed heterogeneity via compressed/dilated time scales.
For completeness, we provide the trajectory-generating equations used in these eight scenarios. Each trajectory includes an additive Gaussian noise term $\epsilon_{ij}\sim N(0,0.8^{2})$. Heterogeneity in progression speed is encoded via a subject-specific factor $s_{i}^{(\cdot)}$ where applicable.
@@ A.2 Scenario Design
Two groups with similar shapes but different means:
$\text{Group 1 (500 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (500 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.$
@@
Two groups with similar means but different shapes:
$\text{Group 1 (500 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=17+2\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (500 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=2t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.$
@@
Two groups with shared shapes and different means; a third group with a different shape:
$\text{Group 1 (300 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=20+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (400 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 3 (300 subjects):}\quad Y_{i}^{(3)}(t_{ij})$ $=16+0.5t_{ij}+3\sin(0.9(t_{ij}+4))+\epsilon_{ij}.$
@@
Four groups with overlapping shapes and means:
$\text{Group 1 (250 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (250 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 3 (250 subjects):}\quad Y_{i}^{(3)}(t_{ij})$ $=38-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij},$
$\text{Group 4 (250 subjects):}\quad Y_{i}^{(4)}(t_{ij})$ $=21+0.9t_{ij}+3\sin(0.6(t_{ij}+4))+\epsilon_{ij}.$
@@
Four groups with two shared shapes and different means:
$\text{Group 1 (250 subjects):}\quad Y_{i}^{(1)}(t_{ij})$ $=24+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 2 (250 subjects):}\quad Y_{i}^{(2)}(t_{ij})$ $=17+3\sin(0.6(t_{ij}+4))+\epsilon_{ij},$
$\text{Group 3 (250 subjects):}\quad Y_{i}^{(3)}(t_{ij})$ $=28-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij},$
$\text{Group 4 (250 subjects):}\quad Y_{i}^{(4)}(t_{ij})$ $=21-0.5t_{ij}+4\sin(0.9(t_{ij}+3))+\epsilon_{ij}.$